# Projecting Assumptions: The Duality Between Sparse Autoencoders and Concept Geometry

### v1

@@ 5.4 Vision
Dataset: We use Imagenette, a 10-class subset of ImageNet, containing 1.5k images per class. Representations are extracted from the DINOv2-base (with registers), yielding 261 tokens per image. Over the course of 50 training epochs, this yields approximately 200 million tokens.
Experiment: SAEs are trained on all available tokens, including spatial, CLS, and registers tokens, for 50 epochs with 200 latent dimensions. For each SAE, the best reconstruction is selected based on a sparsity-controlled learning rate sweep. This resulted in an optimal learning rate of $5\times 10^{-4}$ for TopK, ReLU, and SpaDE, while JumpReLU performed best with $10^{-3}$. Additionally, JumpReLU employs a Silverman kernel with a bandwidth of $10^{-2}$.
Observations: Results are shown in Fig. 8. While a three-dimensional PCA captures a small fraction of total variance ($24\%$), it illustrates the structure of different classes (row (a)). SpaDE again tiles the class structure well. Cosine similarities between pairs of data points using SAE latent representations (first column of each SAE in row (b)) show that all SAEs are able to decorrelate different classes in their latent representations. However, the correlations between SAE latents across all data (second column of each SAE in row (b)) shows co-occurrence across concepts for all SAEs. While this co-occurrence is widespread in ReLU, JumpReLU and TopK SAEs, it seems to be specific to certain pairs of latents in SpaDE. $F_{1}$ scores (row (c)) show that SpaDE has the most monosemantic latents across all classes, even achieving a perfect $F_{1}$ score of 1 for a few classes (Tench, English Springer, French horn and Gas Pump). The varying $F_{1}$ scores for ReLU and JumpReLU across classes indicate different levels of linear separability across classes. Importantly, we find SpaDE identifies interpretable concepts such as foreground/background, different parts of objects in an image (hands, face, fins of fish, windows/ stairs in church images, eyes, ears, snout of dogs, etc), which are visualized using feature attribution maps in App. E.4.
@@ 6 Discussion and Limitations
Our findings reveal critical insights into the limitations and strengths of different sparse autoencoder (SAE) architectures for concept discovery. We observed that ReLU and JumpReLU SAEs fail to capture nonlinear separability, resulting in low F1 scores on such concepts. Additionally, Top-K SAEs struggle to capture concept heterogeneity, performing poorly on concepts with higher intrinsic dimensionality than the chosen value of k. A common issue across these architectures is the co-occurrence of latents across multiple concepts, indicating a lack of concept specialization. In contrast, SpaDE achieves the highest F1 scores for its most monosemantic latents, exhibits low latent co-occurrence across concepts and enforces adaptive sparsity, making it an effective choice for structured concept representations.
These observations highlight that the failure modes of different SAEs stem from a mismatch between their inductive biases and the true structure of the data. Specifically, ReLU and JumpReLU assume linear separability of concepts, which does not always hold, even for concepts that correspond to specific directions in the latent space. This finding is particularly significant, as it challenges the motivation behind ReLU SAEs as following from the linear representation hypothesis (Bricken et al., 2023). On the other hand, our results suggest that incorporating data geometry into SAE design significantly improves concept specialization of SAE latents, allowing it to learn latents that are better aligned with the data.
@@
Limitations:
While SpaDE demonstrates promising improvements over ReLU, JumpReLU and TopK SAEs in synthetic, semi-synthetic and realistic data, we do not claim it to be the optimal SAE for all scenarios. Instead, we present it as a concrete example of how incorporating reasonable data properties— in this case nonlinear separability and concept heterogeneity—can improve interpretability. Thus, several limitations remain, as follows.
  • Data properties beyond those considered here may be crucial for improved SAE performance. Future work should explore additional geometric structure of concepts in neural networks to design better SAEs.
@@
  • We have focused our attention on mutually exclusive concepts in this work, where the presence of one concept implies the absence of others. While our arguments about SAE assumptions hold even when concepts overlap, the expected co-occurrence structure may differ in such cases.
Overall, we note our work is not a proposal for the best SAE, but as a guiding framework for improving design of SAEs that yield useful interpretations, demonstrating how better integration of data geometry can enhance model interpretability. Our findings thus suggest that the interpretability community should prioritize a deeper understanding of latent space geometry. Future research should explore how improved assumptions about data structure can lead to more effective SAE architectures. By systematically incorporating domain-specific insights into SAE design, we can move toward models that provide more faithful and structured representations of concepts.
@@ Acknowledgments
ESL thanks Hidenori Tanaka and the Physics of Intelligence group at CBS-NTT program Harvard for useful conversations and access to compute resources. SSRH thanks the CRISP group at Harvard SEAS for insightful conversations, and the Kempner Institute at Harvard University for access to compute resources.
@@ Appendix A Dictionary Learning
Sparse coding Olshausen & Field (1996a) (alternatively known in this work as sparse dictionary learning, or just dictionary learning) was initially proposed to replicate the observed properties (”spatially localized, oriented, bandpass receptive fields”) of biological neurons in the mammalian visual cortex. It aims to invert a linear generative model with a sparsity prior on the latents:
@@
where $\bm{x}\in\mathbb{R}^{n}$ is the data, $\bm{D}^{*}\in\mathbb{R}^{n\times s}$ is the set of $s$ dictionary atoms, $\bm{z}^{*}\in\mathbb{R}^{s}_{+}$ is the sparse code, and $\eta$ is additive white Gaussian noise. Given data $\{\bm{x}^{(1)},\dots,\bm{x}^{(P)}\}$, sparse coding performs maximum aposteriori (MAP) estimation for the dictionary $\bm{D}^{*}$ and representations $\bm{z}^{*}$ under suitably defined prior and likelihood functions Elad et al. (2007) by solving the following optimization problem (repeated from Eq. 1):
@@
where $\mathcal{R}(\cdot)$ is a sparsity-promoting regularizer. The set $\mathcal{B}\subseteq\mathbb{R}^{n\times s}$ includes restriction to unit norm (typical). Generally, the L1 penalty is used as the regularizer term, i.e., $\mathcal{R}(\bm{z}^{(k)})=\|\bm{z}^{(k)}\|_{1}$, since using the L0 penalty makes the problem NP-hard Tillmann (2015). When the number of dictionary atoms is less than or equal to the dimension of input space, $s\leq n$, this is an undercomplete problem, and the sparse code can be readily obtained using the pseudo-inverse of the dictionary matrix $D$ (provided the dictionary atoms are linearly independent), leading to the solution $\bm{z}=(\bm{D}^{T}\bm{D})^{-1}\bm{D}^{T}\bm{x}$. Note that in this (undercomplete) case, the sparse code is a linear transformation of the input. The more interesting setting involves using an overcomplete dictionary ($s>n$), and was initially studied in Olshausen & Field (1997). Obtaining the sparse code $\bm{z}$ from input data $\bm{x}$ is nontrivial in this case.
In this (overcomplete) setting, a popular approach is using iterative shrinkage and thresholding algorithms (ISTA) Daubechies et al. (2004) and their variants such as FISTA (Fast ISTA) Beck & Teboulle (2009). Modern approaches to this problem use ISTA to design deep residual networks with shared weights and train the network on the sparse coding objective, in a technique called Learned ISTA (LISTA) Gregor & LeCun (2010). Algorithm unrolling Monga et al. (2021) is a generalization of this technique and involves designing interpretable neural networks using iterative algorithms where each layer of the network reflects an iteration of the algorithm. These networks are interpretable since the weights correspond to an underlying process which was used to design the iterative algorithm. Unrolling has widespread applications in signal processing, and is extensively reviewed in Monga et al. (2021).
We also note that sparse coding has been used with algorithm unrolling as a model-based interpretable deep learning technique for a wide range of applications, including image super-resolution Wang et al. (2015), graph signal denoising Chen et al. (2021), mechanical fault diagnosis An et al. (2022), deconvolving neural activity of dopamine neurons in mice Tolooshams et al. (2024). Therefore, assuming a linear generative model of data (Eq. 1) where the dictionary atoms are physically relevant in some application, sparse coding using an unrolled network learns the underlying interpretable dictionary atoms.
@@ Appendix B Related Work
SAEs are a specific instantiation of the broader agenda of dictionary learning tools for concept-level explainability (Kim et al., 2018; olah2020an; Fel, 2025; Faruqui et al., 2015; Subramanian et al., 2018; Arora et al., 2018). A number of SAE architectures have been proposed recently, including ReLU SAE (Bricken et al., 2023), TopK SAE (Gao et al., 2024; Makhzani & Frey, 2013), gated SAE (Rajamanoharan et al., 2024a), JumpReLU SAE (Rajamanoharan et al., 2024b), Batch TopK SAE ((Bussmann et al., 2024)), ProLU SAE ((Taggart, 2024)), and so on. While promising results have been discovered, e.g., latents that respond to concepts of refusal, gender, text script (Bricken et al., 2023; Templeton et al., 2024; Durmus et al., 2024), foreground vs. background concepts (Fel et al., 2023), and concepts of protein structures (Simon & Zou, 2024; Garcia & Ansuini, 2025; Adams et al., 2025), a series of negative results have started to emerge on the limitations of SAEs. For example, Bhalla et al. (2024); Wu et al. (2025a) show a mere prompting baseline can outperform model control compared to SAE or probing based feature ablation baseline. Similar results were observed by Menon et al. (2024) in a narrower formal language setting. Meanwhile, criticizing the underlying linear representation hypothesis that has informed design of earlier SAE architectures (specifically, the vanilla ReLU SAEs), Engels et al. (2024a; b) has shown that SAE features can be multidimensional and nonlinear. Importantly, recent results from Fel et al. (2025); Paulo & Belrose (2025); Kissane et al. (2024) have shown that two SAEs trained on the exact same data, just with a different seed, can yield very different concepts and hence very different interpretations. These results are related to the lack of canonical nature in SAE latents (Leask et al., 2025) This behavior, often called algorithmic instability, makes reliability of SAEs challenging for any practical purposes. More broadly, given the hefty research investment going into the topic, we believe it is warranted that a more formal and theoretical account help solidify the limitations and challenges SAEs (or at least the current paradigm thereof) faces. This can help steer the research in a direction that yields meaningful improvement in SAEs, e.g., in their practical utility. This motivation underscores our work. For a related effort on this front, we highlight the work by Ayonrinde et al. (2024), who contextualize SAEs from a minimum-description length perspective and enable an intuitively solid account of how features may split to overly specialized concepts (e.g., tokens).
@@
Disentangled Representation Learning. As mentioned in Sec. 3, results similar to ours have been reported in the field of disentangled representation learning, wherein one aims to invert a data-generating process to identify the factors of variants (i.e., latent variables) that underlie it. To this end, autoencoders were used as a popular tool, since they offer a method that can (ideally) simultaneously invert the generative process and identify the underlying latents (Higgins et al., 2017). However, Locatello et al. (2019) showed that in fact this problem is rather challenging: unless one designs an autoencoder architecture that bakes-in assumptions about the generative process, i.e., the precise function mapping itself, there are no guarantees the retrieved latents will correspond to the ground-truth ones. This result led to design of several methods focused on exploiting “weak supervision”, i.e., extra information available from data-pairs such as multiple views of an image or temporally consistent video frames, to circumvent the theoretical challenges of disentanglement (Locatello et al., 2020; Gresele et al., 2020; Von Kugelgen et al., 2021). Our contributions are similar in nature to these results on disentanglement, but we (i) specifically focus on the context of SAEs and (ii) provide a more concrete proof that establishes precisely what the inductive biases of popular SAEs are, i.e., what concepts the SAEs are biased towards uncovering. Having established these results, we now believe the next step that the disentanglement community took, i.e., use of weak supervision, would make sense for the SAEs community as well. This can involve exploiting temporal correlations between tokens in a sentence, or the fact that representations across layers do not change much, as in Crosscoders and Transcoders (Lindsey et al., 2024; Dunefsky et al., 2025; Paulo et al., 2025).
@@ C.1 Separability experiment
We construct a synthetic dataset consisting of six isotropic Gaussian clusters in a two-dimensional (2D) space. The cluster centers are arranged such that adjacent clusters are separated by an angular difference of $62\pi/$ 6, with alternate clusters having norms of 1 and 3. Each cluster is sampled from a multivariate normal distribution with a variance of $2^{-5.5}$. The dataset consists of 1 million data points per concept, yielding a total of 6 million samples. Of these, we use 70% (700,000 points) for training.
Our experiments evaluate four sparse autoencoder (SAE) architectures: ReLU SAE, JumpReLU SAE, TopK SAE, and SpaDE. The first three architectures are implemented following their original formulations (in Bricken et al. (2023),Rajamanoharan et al. (2024b),Gao et al. (2024)), with the decoder activations normalized in the forward pass. The SpaDE model follows the same single hidden-layer autoencoder structure but differs in that it utilizes Euclidean distance computations and a SparseMax activation function for the encoder. Across all models, the hidden-layer width is set to 128, and a pre-encoder bias is used in all cases except for SpaDE.
[…]
Regularization parameters are selected such that sparsity levels remain comparable across models. Specifically, the regularization coefficient $\gamma$ is chosen in the range $10^{-6}$ to 1 for ReLU and JumpReLU SAEs, between 4 and 64 (powers of 2) for TopK SAE, and in the range $10^{-6}$ to 1 for SpaDE. Each model applies a different regularization strategy: ReLU SAE uses $L_{1}$ regularization, JumpReLU SAE applies $L_{0}$ regularization with a straight-through estimator (STE) as in Rajamanoharan et al. (2024b), TopK SAE does not use explicit regularization but incorporates an auxiliary loss term as in Gao et al. (2024), with $K_{aux}=k$ (same as the choice of sparsity level $k$ in TopK) with $\gamma_{\text{aux}}=1$ (the scaling for the auxiliary loss term), and SpaDE employs a distance-weighted $L_{1}$ regularization, which comes from Tasissa et al. (2023).
@@
Precision=True PositivesTrue Positives+False $\text{Precision}=\frac{\text{True Positives}}{\text{True Positives}+\text{False Positives}},$ (7)
@@
Recall=True PositivesTrue Positives+False $\text{Recall}=\frac{\text{True Positives}}{\text{True Positives}+\text{False Negatives}}.$ (8)
@@
Stable $\text{Stable Rank}=\frac{\sum\sigma_{i}}{\sigma_{\max}}.$ (10)

### v2

@@ 5.4 Vision
Dataset and Experiment: We use Imagenette, a 10-class subset of ImageNet [50], containing 1.5k images per class. Representations are extracted from the DINOv2-base (with registers), yielding 261 tokens per image. Over the course of 50 training epochs, this yields approximately 200 million tokens. SAEs are trained on all available tokens, including spatial, CLS, and registers tokens, for 50 epochs with 200 latent dimensions.
Observations: Results are shown in Fig. 8. SpaDE again tiles the class structure well in the 3-D PCA (row (a)). Similarities between sparse codes of data (first column of each SAE in row (b)) show that all SAEs are able to decorrelate different classes in their latent representations. Latent co-occurrence (second column of each SAE in row (b)) is widespread in ReLU, JumpReLU and TopK SAEs, but it seems to be specific to certain pairs of latents in SpaDE. $F_{1}$ scores (row (c)) show that SpaDE has the most monosemantic latents across all classes. The varying $F_{1}$ scores for ReLU and JumpReLU across classes indicate different levels of linear separability across classes. Importantly, we find SpaDE identifies interpretable concepts such as foreground/background, different parts of objects in an image (hands, face, fins of fish, windows/ stairs in church images, eyes, ears, snout of dogs, etc), which are visualized using feature attribution maps in App. E.4.
@@ 6 Discussion and Limitations
Our findings reveal critical insights into the limitations and strengths of different sparse autoencoder (SAE) architectures for concept discovery. We observed that ReLU and JumpReLU SAEs fail to capture nonlinear separability ( low F1 scores, latent co-occurrence across concepts), while Top-K SAE struggles to capture concept heterogeneity (high MSE when concept dimension exceeds choice of K). A common issue across these architectures is the co-occurrence of latents across multiple concepts, indicating a lack of concept specialization. In contrast, SpaDE achieves the highest F1 scores for its most monosemantic latents, exhibits low latent co-occurrence across concepts and enforces adaptive sparsity, making it an effective choice for structured concept representations. Our observations about the limitations of ReLU, JumpReLU and TopK SAEs highlight that the failure modes of different SAEs stem from a mismatch between their inductive biases and the true structure of the data. Specifically, ReLU and JumpReLU assume linear separability of concepts, which does not always hold, even for concepts that correspond to specific directions in the latent space. On the other hand, our results suggest that incorporating data geometry into SAE design significantly improves concept specialization of SAE latents, allowing it to learn latents that are better aligned with the data.
@@
Our analysis of receptive fields of SAE encoders and their relation with concept geometry to study monosemanticity properties is quite general, and can also be used to study other kinds of interpreter models such as transcoders [51, 52].
Limitations: While SpaDE demonstrates promising improvements over ReLU, JumpReLU and TopK SAEs in synthetic, semi-synthetic and realistic data, we do not claim it to be the optimal SAE for all scenarios. Instead, we present it as a concrete example of how incorporating reasonable data properties (nonlinear separability and concept heterogeneity) can improve interpretability. Thus, several limitations remain, as follows.
  • Data properties beyond those considered here may be crucial for improved SAE performance. Future work may explore additional geometric structure of concepts in neural networks to design better SAEs.
@@
  • We have focused our attention on mutually exclusive concepts, where the presence of one concept implies the absence of others. While our arguments about SAE assumptions hold even when concepts overlap, the expected co-occurrence structure may differ in such cases. For co-occurring concepts, our receptive field analysis can be applied to the presence/absence of each concept.
Overall, we note our work is not a proposal for the best SAE, but a guiding framework for improving design of SAEs that yield useful interpretations demonstrating how better integration of data geometry can enhance model interpretability. The interpretability community may need to prioritize a deeper understanding of latent space geometry, and translate novel insights into SAE design, leading to models with more faithful and structured representations of concepts.
@@ Acknowledgments
This work has been made possible in part by a gift from the Chan Zuckerberg Initiative Foundation to establish the Kempner Institute for the Study of Natural and Artificial Intelligence at Harvard University. ESL thanks Hidenori Tanaka and the Physics of Intelligence group at CBS-NTT program Harvard for useful conversations and access to compute resources. All authors thank the CRISP group at Harvard SEAS for insightful conversations.
@@ Appendix A Dictionary Learning
Sparse coding [10] (alternatively known in this work as sparse dictionary learning, or just dictionary learning) was initially proposed to replicate the observed properties ("spatially localized, oriented, bandpass receptive fields") of biological neurons in the mammalian visual cortex. It aims to invert a linear generative model with a sparsity prior on the latents:
@@
where $\bm{x}\in\mathbb{R}^{n}$ is the data, $\bm{D}^{*}\in\mathbb{R}^{n\times s}$ is the set of $s$ dictionary atoms, $\bm{z}^{*}\in\mathbb{R}^{s}_{+}$ is the sparse code, and $\eta$ is additive white Gaussian noise. Given data $\{\bm{x}^{(1)},\dots,\bm{x}^{(P)}\}$, sparse coding performs maximum aposteriori (MAP) estimation for the dictionary $\bm{D}^{*}$ and representations $\bm{z}^{*}$ under suitably defined prior and likelihood functions [53] by solving the following optimization problem (repeated from Eq. 1):
@@
where $\mathcal{R}(\cdot)$ is a sparsity-promoting regularizer. The set $\mathcal{B}\subseteq\mathbb{R}^{n\times s}$ includes restriction to unit norm (typical). Generally, the L1 penalty is used as the regularizer term, i.e., $\mathcal{R}(\bm{z}^{(k)})=\|\bm{z}^{(k)}\|_{1}$, since using the L0 penalty makes the problem NP-hard [54]. When the number of dictionary atoms is less than or equal to the dimension of input space, $s\leq n$, this is an undercomplete problem, and the sparse code can be readily obtained using the pseudo-inverse of the dictionary matrix $D$ (provided the dictionary atoms are linearly independent), leading to the solution $\bm{z}=(\bm{D}^{T}\bm{D})^{-1}\bm{D}^{T}\bm{x}$. Note that in this (undercomplete) case, the sparse code is a linear transformation of the input. The more interesting setting involves using an overcomplete dictionary ($s>n$), and was initially studied in [35]. Obtaining the sparse code $\bm{z}$ from input data $\bm{x}$ is nontrivial in this case.
In this case, sparse coding results in a sparse representation of the data and a dictionary which behaves as a data-adaptive basis. Correspondingly, sparse codes have been shown to capture interesting concepts in data [55, 56], e.g., responding to wavelet-like regions when trained on natural images [57]. In this (overcomplete) setting, a popular approach is using iterative shrinkage and thresholding algorithms (ISTA) [58] and their variants such as FISTA (Fast ISTA) [59]. Modern approaches to this problem use ISTA to design deep residual networks with shared weights and train the network on the sparse coding objective, in a technique called Learned ISTA (LISTA) [60]. Algorithm unrolling [61] is a generalization of this technique and involves designing interpretable neural networks using iterative algorithms where each layer of the network reflects an iteration of the algorithm. These networks are interpretable since the weights correspond to an underlying process which was used to design the iterative algorithm. Unrolling has widespread applications in signal processing, and is extensively reviewed in [61].
We also note that sparse coding has been used with algorithm unrolling as a model-based interpretable deep learning technique for a wide range of applications, including image super-resolution [62], graph signal denoising [63], mechanical fault diagnosis [64], deconvolving neural activity of dopamine neurons in mice [65]. Therefore, assuming a linear generative model of data (Eq. 1) where the dictionary atoms are physically relevant in some application, sparse coding using an unrolled network learns the underlying interpretable dictionary atoms.
@@ Appendix B Related Work
SAEs are a specific instantiation of the broader agenda of dictionary learning tools for concept-level explainability [19, 66, 20, 7, 8, 9]. A number of SAE architectures have been proposed recently, including ReLU SAE [12], TopK SAE [13, 37], gated SAE [67], JumpReLU SAE [14], Batch TopK SAE ([16]), ProLU SAE ([68]), and so on. While promising results have been discovered, e.g., latents that respond to concepts of refusal, gender, text script [12, 22, 23], foreground vs. background concepts [17], and concepts of protein structures [25, 27, 26], a series of negative results have started to emerge on the limitations of SAEs. For example, [33, 69] show a mere prompting baseline can outperform model control compared to SAE or probing based feature ablation baseline. Similar results were observed by [34] in a narrower formal language setting. Meanwhile, criticizing the underlying linear representation hypothesis that has informed design of earlier SAE architectures (specifically, the vanilla ReLU SAEs), [42, 70] has shown that SAE features can be multidimensional and nonlinear. Importantly, recent results from [15, 32, 71] have shown that two SAEs trained on the exact same data, just with a different seed, can yield very different concepts and hence very different interpretations. These results are related to the lack of canonical nature in SAE latents [72] This behavior, often called algorithmic instability, makes reliability of SAEs challenging for any practical purposes. More broadly, given the hefty research investment going into the topic, we believe it is warranted that a more formal and theoretical account help solidify the limitations and challenges SAEs (or at least the current paradigm thereof) faces. This can help steer the research in a direction that yields meaningful improvement in SAEs, e.g., in their practical utility. This motivation underscores our work. For a related effort on this front, we highlight the work by [73], who contextualize SAEs from a minimum-description length perspective and enable an intuitively solid account of how features may split to overly specialized concepts (e.g., tokens).
@@
Disentangled Representation Learning. As mentioned in Sec. 3, results similar to ours have been reported in the field of disentangled representation learning, wherein one aims to invert a data-generating process to identify the factors of variants (i.e., latent variables) that underlie it. To this end, autoencoders were used as a popular tool, since they offer a method that can (ideally) simultaneously invert the generative process and identify the underlying latents [74]. However, [75] showed that in fact this problem is rather challenging: unless one designs an autoencoder architecture that bakes-in assumptions about the generative process, i.e., the precise function mapping itself, there are no guarantees the retrieved latents will correspond to the ground-truth ones. This result led to design of several methods focused on exploiting “weak supervision”, i.e., extra information available from data-pairs such as multiple views of an image or temporally consistent video frames, to circumvent the theoretical challenges of disentanglement [76, 77, 78]. Our contributions are similar in nature to these results on disentanglement, but we (i) specifically focus on the context of SAEs and (ii) provide a more concrete proof that establishes precisely what the inductive biases of popular SAEs are, i.e., what concepts the SAEs are biased towards uncovering. Having established these results, we now believe the next step that the disentanglement community took, i.e., use of weak supervision, would make sense for the SAEs community as well. This can involve exploiting temporal correlations between tokens in a sentence, or the fact that representations across layers do not change much, as in Crosscoders and Transcoders [79, 51, 52].
@@ Appendix C Experimental Setup
The synthetic experiments (separability, heterogeneity) and vision experiments were run on NVIDIA A100 40GB GPUs, while the formal language experiments were run on NVIDIA RTX A6000 48GB GPUs.
@@ C.1 Separability experiment
We construct a synthetic dataset consisting of six isotropic Gaussian clusters in a two-dimensional (2D) space. The cluster centers are arranged such that adjacent clusters are separated by an angular difference of $2\pi/6$, with alternate clusters having norms of 1 and 3. Each cluster is sampled from a multivariate normal distribution with a variance of $2^{-5.5}$. The dataset consists of 1 million data points per concept, yielding a total of 6 million samples. Of these, we use 70% (700,000 points) for training.
Our experiments evaluate four sparse autoencoder (SAE) architectures: ReLU SAE, JumpReLU SAE, TopK SAE, and SpaDE. The first three architectures are implemented following their original formulations (in [12],[14],[13]), with the decoder activations normalized in the forward pass. The SpaDE model follows the same single hidden-layer autoencoder structure but differs in that it utilizes Euclidean distance computations and a SparseMax activation function for the encoder. Across all models, the hidden-layer width is set to 128, and a pre-encoder bias is used in all cases except for SpaDE.
For training, the (inverse) temperature parameter $\lambda$ in SpaDE is initialized to $1/(2\times\text{input dimension})$ and parameterized using the Softplus function to ensure non-negativity. This parameter trained along with the encoder and decoder weights, to allow the model to learn its desired sparsity level. Note that large values of $\lambda$ lead to greater sparsity since Sparsemax is scale-sensitive. In JumpReLU, the threshold is initialized at $10^{-3}$ across all latent dimensions, with a bandwidth of $10^{-3}$ for the straight-through estimator (STE), as it is proposed in [14]. All models are trained using the Adam optimizer with a learning rate of $10^{-2}$, which follows a cosine decay schedule from $10^{-2}$ to $10^{-4}$. The momentum parameter is set to 0.9, and we use a batch size of 512. Training runs for approximately 8000 iterations, and gradient clipping is applied (gradient norms are clipped at 1) to stabilize optimization.
Regularization parameters are selected such that sparsity levels remain comparable across models. Specifically, the regularization coefficient $\gamma$ is chosen in the range $10^{-6}$ to 1 for ReLU and JumpReLU SAEs, between 4 and 64 (powers of 2) for TopK SAE, and in the range $10^{-6}$ to 1 for SpaDE. Each model applies a different regularization strategy: ReLU SAE uses $L_{1}$ regularization, JumpReLU SAE applies $L_{0}$ regularization with a straight-through estimator (STE) as in [14], TopK SAE does not use explicit regularization but incorporates an auxiliary loss term as in [13], with $K_{aux}=k$ (same as the choice of sparsity level $k$ in TopK) with $\gamma_{\text{aux}}=1$ (the scaling for the auxiliary loss term), and SpaDE employs a distance-weighted $L_{1}$ regularization, which comes from [45].
@@
$\text{Precision}=\frac{\text{True Positives}}{\text{True Positives}+\text{False Positives}},$ (7)
@@
$\text{Recall}=\frac{\text{True Positives}}{\text{True Positives}+\text{False Negatives}}.$ (8)
@@
$\text{Stable Rank}=\frac{\sum\sigma_{i}}{\sigma_{\max}}.$ (10)